OpenStudy (anonymous):

Twelve is the geometric mean of 9 and what number?

OpenStudy (amistre64):

a geometric mean of 2 numbers is the sqrt of the product

OpenStudy (amistre64):

\[12 = \sqrt{9n}\]

OpenStudy (anonymous):

because the geometric mean can be given by \[\sqrt{b*a}\] for 2 numbers, b and a, \[12 = \sqrt{9*x}\] solve for x

OpenStudy (anonymous):

am i suppose to divide?

OpenStudy (anonymous):

is the answer 4?

OpenStudy (amistre64):

youre spose to work the algebra ... square then divide

OpenStudy (amistre64):

and no, 12 is not equal to the sqrt(36)

OpenStudy (anonymous):

24?

OpenStudy (amistre64):

no ... are you guessing? we cant see where you are going astray

OpenStudy (anonymous):

no imnot im not good at geometry at all

OpenStudy (amistre64):

your in luck then, this isnt geometry

OpenStudy (anonymous):

im sorry if im waisting ur time here let me try again and im in a geometry coarse haha :p but ok

OpenStudy (amistre64):

its not a wasting time :)\[12=\sqrt{9n}\] square both side\[12^2=[\sqrt{9n}]^2\] \[144=9n\] now divide the 9\[n=144/9\]

OpenStudy (anonymous):

o my god i get it it knowwwwww thank u so much the answers 16

OpenStudy (amistre64):

geometry-wise its the proportion of similar triangles:\[\frac{12}{9}=\frac{n}{12}\]

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thank you :)